# Glassy dynamics in deep neural networks: A structural comparison

### v2

@@ I Introduction
Deep neural networks (DNNs) have experienced an explosion of popularity in the last decade as they have proven to be highly effective at tackling a wide range of problems across the natural sciences and engineering. Consequently, there is significant interest from the statistical physics community in understanding the fundamental mechanisms by which neural networks train and generalise [1, 2, 3, 4, 5]. An important strand of this work is the similarity between DNNs and glasses. The link between spin glasses and neural networks has received considerable attention since the 1980s, with an emphasis on the mapping between the Hopfield network and Ising-like systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. More recently, the authors of [18] identified both points of similarity and difference between the mean square displacement of a spin glass and a variety of feed forward networks, trained with Stochastic Gradient Descent (SGD). However, the possible analogy between neural networks and structural glasses, which exhibit a richer phenomenology than spin glasses, is still largely unexplored. Moreover, since modern deep learning comprises a wide variety of (typically feed forward, as opposed to Hopfield) architectures, one may also wonder whether this analogy might be extended to feed forward DNNs with more complex interactions between parameters.
@@
Structural glasses exhibit a wide variety of typical phenomena (growing and ultimately diverging relaxation times, caging, dynamic heterogeneity, violation of the Stokes-Einstein relation, ageing) that have been the subject of intense research for many decades, resulting in a rich body of theory of the kind that is currently lacking in deep learning [19, 20, 21, 22]. Structural glasses are a potentially valuable analogy for DNNs as they involve complicated interaction potentials between the constituent particles. This is reminiscent of the situation in DNNs where parameters also interact via highly non-linear, indirect couplings. Furthermore, recent studies have made a quantitative comparison between the jamming transition (a phenomenon closely related to the structural glass transition [23]) and an under-to-overparameterised transition (UOT) in DNNs [24, 25, 26, 27]. Large networks are able to fit their training data exactly, resulting in zero loss (equivalent to an unjammed state with zero interaction energy), whereas small networks have a non-zero loss similar to the positive interaction energy of jammed particles. Much like the energy landscape of an unjammed system, the loss landscape in the long-time limit of overparameterised networks has been shown to consist of a large number of flat directions [18]. These recent results raise the intriguing possibility, explored in this work, of whether similar connections can be drawn between the dynamics of glass forming particle systems and DNNs.
@@
As well as the analogy with jamming, results derived from high-dimensional random energy landscapes [28] are also relevant to the statistical physics of deep learning. Random energy landscapes, which are in some respects highly idealised models of glasses, have been widely applied in physics [29, 30, 31] and are now increasingly being used to study DNNs and related machine learning systems [32, 33, 34]. Of particular relevance to the current work is the Topology Trivialisation Transition (TTT), where the energy landscape of a system switches from a single minimum to many minima on decreasing a quadratic energy penalty. The existence of TTT-like transitions has been demonstrated in high-dimensional elastic models subject to random fields [35, 36], in large ecosystems [37], and in random neural networks [32, 13]. These diverse systems all share a common structure with feed forward DNNs subject to regularisation in that their dynamics balance a disordered term with a quadratic potential. It has remained unexplored, to the best of the authors’ knowledge, whether regularised feed forward DNNs undergo a TTT (and in fact, we find that they do.)
@@ II Numerical setup
In this work we use fully connected feedforward networks with no biases, only weights, so that all parameters are of the same type. The DNNs are denoted $f(\mathbf{x},\mathbf{W})$ for weight vector $\mathbf{W}$, of length $N$, and input data $\mathbf{x}$. The activation functions are ReLU, $\sigma(z)=\text{max}(0,z)$. We consider a problem of supervised binary classification, with a training set $\mathcal{T}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{P_{D}}$, of size $P_{D}$, consisting of input examples $\mathbf{x}$ of dimension $d$, and corresponding labels $y\in\{+1,-1\}$.
@@
where $\lambda>0$ is a control parameter. When studying dynamic variables, we train the DNNs with overdamped Langevin dynamics (also known as Stochastic Gradient Langevin Dynamics, SGLD, in the deep learning literature),
@@
where $w_{i}$ is a weight, $\mu$ is the learning rate (set to $10^{-3}$ unless otherwise stated), and the noise term obeys $\langle\xi_{i}\rangle=0$, and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=(2k_{B}T\mu/P_{D})\delta_{ij}\delta(t-t^{\prime})$, for temperature $T$ and Boltzmann constant $k_{B}$. For the remainder of the paper we work in natural units where $k_{B}=1$. In contrast to physical systems a factor of $1/P_{D}$, the size of the dataset, appears in the noise correlation due to $\mathcal{L}$ being intensive with respect to the dataset size. See Appendix A for a more detailed discussion. Equation (2) highlights the similarity between DNNs and the disordered elastic model of [28], which undergoes a TTT. The Hamiltonian of the elastic system is $\mathcal{H}(x_{i})=V(x_{i})+\frac{\eta}{2}|\mathbf{x}|^{2}$, where $V$ is a random potential, and the second term describes the elasticity of the medium. This is also similar to the Hamiltonian of spherical spin glasses [41]. The question we wish to address is, in essence, how much glassy behaviour remains when we substitute $\mathcal{H}$ for $\mathcal{L}$.
We will now consider various aspects of glassy dynamics, each of which requires a different protocol to potentially probe it in DNNs. We will discuss both the protocol and the results, starting with the phase diagram in order to identify the transition lines. We will then address relaxation dynamics in the resulting disordered phases, time-temperature superposition, caging and the mean square displacement, the Stokes-Einstein relation, non-Gaussianity, dynamic heterogeneity, and finally end with aging. Concluding comments are given in Sec. IV.
@@ III.1 Phase diagram
We first seek to establish a phase diagram to identify two relevant transition lines for our DNNs. The first is a transition in the DNN loss as a function of the ratio of network size, $N$, to dataset size, $P_{D}$, from underparameterised ($\mathcal{L}>0$) to overparameterised ($\mathcal{L}\approx 0$). This transition is analogous to the jamming transition of interacting particles, and was discovered by the authors of [24]. The second transition, the Topology Trivialisation Transition of [28], separates a regime with a single energy minimum, from a disordered energy landscape with many minima. As discussed below, we find the same under-to-overparameterised/jamming transition, and show that an analogous TTT exists in DNNs as a function of the regularisation strength, $\lambda$.
@@
In order to construct the phase diagram, we train DNNs with Stochastic Gradient Descent in the $(N,\lambda)$ plane, for 7000 epochs with a batch size of 100. We then perform 50 epochs of gradient descent at the original learning rate, and a further 50 epochs of gradient descent after reducing the learning rate by a factor of 10. Convergence was checked manually on a representative subset of DNNs. As such we are sampling (close to) the minima of the landscape. Finally we average over 10 DNNs with different, random initial conditions at each point in the $(N,\lambda)$ plane, resulting in the phase diagram shown in Fig. 1(a). We use SGD instead of Langevin dynamics as we are studying static properties of the landscape, not the dynamics, and SGD is much faster than Eq. (3).
@@
The UOT is visible on the $\lambda=0$ line, whereas the TTT occurs in the vertical direction. The transitions are shown separately in Fig. 1(b) (UOT) and (c) (TTT). Figure 1(b) uses the same networks as on the $\lambda=0$ line from Fig. 1(a), and shows that achieving the $\mathcal{L}\approx 0$ state is by no means guaranteed. It is strongly dependent on the initial conditions and training trajectory through weight space, even in the overparameterised phase, in agreement with [24]. On either side of the UOT we have two disordered phases, underparameterised, and overparameterised, illustrated schematically in Fig. 1(a). The key difference between these phases is the presence of flat plateaus at zero loss in the overparameterised phase. The UOT does not exist when $\lambda>0$ as the addition of the regularisation term destroys the plateaus. This is similar to the case where the addition of a spherical constraint makes a system of equations unsolvable [42]. Figure 1(c) uses 20 DNNs of width 30, each with different initial conditions, to demonstrate the TTT. Each DNN undergoes a sharp TTT, but the position of the transition varies, resulting in an average with some width. In this transition region a small number of initial conditions find a long-lived, high loss minimum, resulting in a kink in the average curve. The TTT separates a trivial phase, where the landscape is essentially a quadratic bowl (high $\lambda$), from a disordered phase (low $\lambda$).
It is possible to make a heuristic argument about how regularisation affects the number of minima in the loss landscape. Turning points in $\mathcal{L}$ can be classified by the eigenvalues of the Hessian, $H_{ij}=\partial^{2}\mathcal{L}/\partial w_{i}\partial w_{j}$, for weights $w_{i}$ and $w_{j}$. A turning point with only positive eigenvalues is a minimum, only negative eigenvalues a maximum, and anything in between a saddle point. $H$ is essentially a random matrix, with some distribution of eigenvalues. Adding L2 regularisation, as in Eq. (2), transforms the Hessian to $H_{ij}=\partial^{2}\mathcal{L}_{\text{data}}/\partial w_{i}\partial w_{j}+\lambda\delta_{ij}$, with corresponding characteristic polynomial, $\text{det}(H-(\sigma-\lambda)I)=0$, for eigenvalues $\sigma$, and identity matrix, $I$. Hence the regularisation term, $\lambda$, shifts the eigenvalue spectrum towards positive values, and makes minima more likely.
@@
Figure 1: Glass-like transitions in neural networks. (a) The final training loss (colour bar) in the $(N/P_{D},\lambda)$ plane, where $N$ is the number of weights, $P_{D}$ the number of training data, and $\lambda$ the regularisation strength. Schematic illustrations of the loss landscape in the resulting four phases are shown around the edge. At high $\lambda$ the quadratic L2 regularisation term of Eq. (2) dominates and the loss is a simple well. As $\lambda$ decreases the loss landscape undergoes a TTT and develops many random minima. On the $\lambda=0$ line, at low $N/P_{D}$ the network is underparameterised, corresponding to another disordered landscape. Finally, at high $N/P_{D}$ the network is over parameterised and has a loss landscape combining disordered regions with flat plateaus at zero loss. (b) The final training loss either side of the UOT. Each black point is a different initial condition. (c) The final training loss on either side of the TTT. The translucent lines correspond to different initial conditions, and the blue line is their average.
@@ III.2 Relaxation with overdamped Langevin dynamics
We now turn to the time-dependent dynamics of our DNNs to explore potential analogies with the relaxation dynamics of structural glassformers. There are four regions of interest in our phase diagram, on either side of each transition: the regularised phase ($0<\lambda<\lambda^{\ast}$); the quadratic phase ($\lambda>\lambda^{\ast}$); the underparameterised phase ($\lambda=0$, $N<N^{\ast}$); and the overparameterised phase ($\lambda=0$, $N>N^{\ast}$). In practice, the quadratic phase is trivial as networks rapidly reach steady state at the bottom of the quadratic well ($w_{i}\approx 0$ $\forall i$), hence we do not discuss this phase further. As can be seen from Fig. 1(a), $\lambda^{\ast}\approx 0.03$, with some $N$ dependence, and $N^{\ast}\approx 0.18$.
@@
We perform a set of numerical experiments designed to be similar to the quench and anneal protocol used to study glasses. Networks are first initialised with random weights drawn from a uniform distribution. Next, the networks are trained with SGD, at batch size 100, to efficiently move them away from their random initial position in the loss landscape, and into a local basin. The networks are then trained for a further period using the overdamped Langevin dynamics of Eq. (3), at a range of temperatures, $T$. After waiting some time, $t_{w}$, to allow the networks to move beyond their transient response to the new dynamics we start to measure $Q$. Measurements are made on approximately 50 independent initial conditions (some networks are lost as their training dynamics diverge). This protocol is illustrated schematically in Fig. 2(a), and a typical example is shown in Fig. A2. The decay of $Q$ at two temperatures is shown in Fig. 2(b). As expected, networks decorrelate more quickly at higher temperatures. We take as a typical decorrelation time $\tau_{90}$, the time it takes for $Q$ to reach 90% of its initial value, $Q(\tau_{90})=0.9$. We choose $\tau_{90}$ instead of the more typical e-fold relaxation time, $\tau_{e}$, which probes the long time ($\alpha$-relaxation) behaviour, as we do not observe a two stage relaxation process in any of our experiments (e.g. see the left hand column of Fig. 3). Hence any choice of threshold is somewhat arbitrary, and we use 90% for computational convenience as it is reached quickly. For completeness we repeat our measurements with $\tau_{e}$ and find no significant differences with $\tau_{90}$. These results are presented in Appendix E.
@@
Figure 2: (a) A schematic of the measurement procedure. Networks are generated with random weights, and trained for an initial period with SGD. The networks are then trained for a second period with overdamped Langevin dynamics. Measurements start after a waiting time, $t_{w}$. (b) Examples of the overlap function for networks in the high temperature, noise dominated regime (red, lower curve, $T=10^{-4}$), and in the low temperature, landscape dominated regime (blue, upper curve, $T=10^{-6}$). These measurements were made with $t_{w}=5\times 10^{3}$, and using $\chi=400$. The analytic predictions of $Q(t)$ for free Brownian particles are shown by the dashed black lines.
@@
In Fig. 3(a), (b), and (c), $Q(t)$ is plotted for a range of temperatures in the regularised (width=30, $\lambda=10^{-2}$), underparameterised (width=30, $\lambda=0$), and overparameterised (width=72, $\lambda=0$) phases respectively. The corresponding dependence of $\tau_{90}$ on the temperature, $T$, is shown in (d), (e), and (f). Both the regularised and underparameterised phases display two dynamical regimes. At high $T$, decorrelation is dominated by thermal noise, and we find a power law relationship, $\tau_{90}\sim T^{-1}$, which we call the noise-dominated regime. This is what we would expect from an overdamped, memoryless system with dynamics of the form $\dot{Q}=-\Omega Q$, where the frequency term $\Omega\sim T$ (e.g. Eq. 28 in [43] if we neglect the memory kernel). Interestingly, a $1/T$ relationship has also been observed for the decorrelation of the loss of DNNs trained on random data, and using the cross entropy loss function [44]. At low $T$, the decorrelation process becomes temperature independent, and instead is dominated by the first term of Eq. (3), the deterministic gradients of the loss landscape. The mean value of $\tau_{90}$ at $T=0$ is shown with the horizontal black line. Consequently, the value $T^{*}$ where plots (d), (e), and (f) exhibit a kink is the point at which the networks become trapped in the basins of the loss landscape. We call this low $T$ behaviour the landscape-dominated regime. In the overparameterised phase only the noise-dominated regime, $\tau_{90}\sim T^{-1}$, exists. We suggest that this is because the landscape consists of flat plateaus at $\mathcal{L}\approx 0$ where there are no gradients and hence the noise term of Eq. (3) always dominates. Furthermore, the observed divergence of $\tau_{90}$ as $T\to 0$ is to be expected, as both terms in Eq. (3) disappear.
@@
where $\text{erf}(x)$ is the error function. This analytic curve is shown in the dashed line of Fig. 2(b) where, for a network in the noise-dominated regime, the prediction closely follows the DNN measurement at short and intermediate times. In the landscape-dominated regime, the DNNs decorrelate faster than predicted. Weights in a harmonic basin subjected to thermal noise decorrelate faster the steeper the basin (see Appendix F for details), hence, although we cannot perform analytic calculations with the true loss landscape, it is to be expected that DNNs in the landscape-dominated regime decorrelate faster than the prediction of Eq. (6).

### v3

# Do deep neural networks behave like structural glasses?

@@ I Introduction
Deep neural networks (DNNs) have experienced an explosion of popularity in the last decade as they have proven to be highly effective at tackling a wide range of problems across the natural sciences and engineering. Consequently, there is significant interest from the statistical physics community in understanding the fundamental mechanisms by which neural networks train and generalise [1, 2, 3, 4, 5]. An important strand of this work is the similarity between DNNs and glasses. The link between spin glasses and neural networks has received considerable attention since the 1980s, originally with an emphasis on the mapping between the Hopfield network and Ising-like systems, with attention turning to feed-forward networks in the last decade or so [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Several studies have focused on the nature of the loss landscape, in particular demonstrating the surprising result that in many cases the landscape, while non-convex, has no poorly performing local minima [20, 21, 22]. More recently, training dynamics of networks trained with Stochastic Gradient Descent (SGD) have been investigated in the context of anomalous diffusion [23], and in comparison with spin glasses where both similarities and differences were found [24]. There is also some significant overlap between spin and structural glass phenomenology [25], particularly in the limit of infinite dimensions [26, 27, 28, 29]. However, the possible analogy between neural networks and structural glasses, which in general exhibit a richer phenomenology than spin glasses, is still largely unexplored. Moreover, since modern deep learning comprises a wide variety of (typically feed forward, as opposed to Hopfield) architectures, one may also wonder whether this analogy might be extended to feed forward DNNs with more complex interactions between parameters.
@@
Structural glasses exhibit a wide variety of typical phenomena (growing and ultimately diverging relaxation times, caging, dynamic heterogeneity, violation of the Stokes-Einstein relation, ageing) that have been the subject of intense research for many decades, resulting in a rich body of theory of the kind that is currently lacking in deep learning [30, 31, 32, 33]. Structural glasses are a potentially valuable analogy for DNNs as they involve complicated interaction potentials between the constituent particles. This is reminiscent of the situation in DNNs where parameters also interact via highly non-linear, indirect couplings. Furthermore, recent studies have made a quantitative comparison between the jamming transition (a phenomenon closely related to the structural glass transition [34]) and an under-to-overparameterised transition (UOT) in DNNs [35, 36, 37, 38]. Large networks are able to fit their training data exactly, resulting in zero loss (equivalent to an unjammed state with zero interaction energy, and sometimes referred to as a canyon landscape [39, 40]), whereas small networks have a non-zero loss similar to the positive interaction energy of jammed particles. It is highly counter-intuitive that modern deep learning works so successfully in this overparameterised phase while avoiding the problems of overfitting. Much like the energy landscape of an unjammed system, the loss landscape in the long-time limit of overparameterised networks has been shown to consist of a large number of flat directions [24]. These recent results raise the intriguing possibility, explored in this work, of whether similar connections can be drawn between the dynamics of glass forming particle systems and DNNs.
@@
As well as the analogy with jamming, results derived from high-dimensional random energy landscapes [41] are also relevant to the statistical physics of deep learning. Random energy landscapes, which are in some respects highly idealised models of glasses, have been widely applied in physics [42, 43, 44] and are now increasingly being used to study DNNs and related machine learning systems [45, 46, 47]. Of particular relevance to the current work is the Topology Trivialisation Transition (TTT), where the energy landscape of a system switches from a single minimum to many minima on decreasing a quadratic energy penalty. The existence of TTT-like transitions has been demonstrated in high-dimensional elastic models subject to random fields [48, 49], in large ecosystems [50], in systems of coupled differential equations [51], and in random neural networks [45, 13]. These diverse systems all share a common structure with feed forward DNNs subject to regularisation in that their dynamics balance a disordered term with a quadratic potential. It has remained unexplored, to the best of the authors’ knowledge, whether regularised feed forward DNNs undergo a TTT in their loss function (and in fact, we find that they do.)
@@ II Numerical setup
In this work we use fully connected feedforward networks with no biases, only weights, so that all parameters are of the same type. We use this structure as it is the most general feed forward architecture, while noting that our DNNs differ considerably both in structure and size from common models such as convolutional neural networks and transformers. The output of a DNN is denoted $f(\mathbf{x},\mathbf{W})$ for weight vector $\mathbf{W}$, of length $N$, and input data $\mathbf{x}$. The activation functions are ReLU, $\sigma(z)=\text{max}(0,z)$. We consider a problem of supervised binary classification, with a training set $\mathcal{T}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{P_{D}}$, of size $P_{D}$, consisting of input examples $\mathbf{x}$ of dimension $d$, and corresponding labels $y\in\{+1,-1\}$.
@@
where $\lambda>0$ is a control parameter. When studying dynamic variables, we train the DNNs with overdamped Langevin dynamics,
@@
where $w_{i}$ is a weight, $\mu$ is the learning rate (set to $10^{-3}$ unless otherwise stated), and the noise term obeys $\langle\xi_{i}\rangle=0$, and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=(2k_{B}T\mu/P_{D})\delta_{ij}\delta(t-t^{\prime})$, for temperature $T$ and Boltzmann constant $k_{B}$. For the remainder of the paper we work in natural units where $k_{B}=1$. In contrast to physical systems a factor of $1/P_{D}$, the size of the dataset, appears in the noise correlation due to $\mathcal{L}$ being intensive with respect to the dataset size. See Appendix A for a more detailed discussion.
At times we also train networks with SGD, where the dynamics are given by
$\partial_{t}w_{i}=-\mu\frac{\partial\tilde{\mathcal{L}}}{\partial w_{i}}.$ (4)
Here the loss is approximated by using only a subset of the data (a “batch”) of size $B$, ~$\tilde{\mathcal{L}}=1/B\sum_{i=1}^{B}\ell(f(\mathbf{x}_{i},\mathbf{W}),y_{i})$. Although no noise appears explicitly in Eq. (4), stochasticity is introduced by the random selection of batches. We do not use SGD for dynamic measurements as the noise is highly unphysical - it is coloured, non-delta correlated, and non-constant in time.
Equation (2) highlights the similarity between DNNs and the disordered elastic model of [41], which undergoes a TTT. The Hamiltonian of the elastic system is $\mathcal{H}(x_{i})=V(x_{i})+\frac{\eta}{2}|\mathbf{x}|^{2}$, where $V$ is a random potential, and the second term describes the elasticity of the medium. This is also similar to the Hamiltonian of spherical spin glasses [55]. The question we wish to address is, in essence, how much glassy behaviour remains when we substitute $\mathcal{H}$ for $\mathcal{L}$.
We will now consider various aspects of glassy dynamics, each of which requires a different protocol to potentially probe it in DNNs. We will discuss both the protocol and the results, starting with the phase diagram in order to identify the transition lines. We will then address relaxation dynamics in the resulting disordered phases, time-temperature superposition, caging and the mean square displacement, the Stokes-Einstein relation, non-Gaussianity, dynamic heterogeneity, and finally end with aging. Concluding comments are given in Sec. IV. The code used to generate and analyse the DNNs is available at https://github.com/mkerrwinter/glass_net_publication.
@@ III.1 Phase diagram
We first seek to establish a phase diagram to identify two relevant transition lines for our DNNs. The first is a transition in the DNN loss as a function of the ratio of network size, $N$, to dataset size, $P_{D}$, from underparameterised ($\mathcal{L}>0$) to overparameterised ($\mathcal{L}\approx 0$). This transition is analogous to the jamming transition of interacting particles, and was first applied to the perceptron in [38], and then to DNNs in [35]. The second transition, the Topology Trivialisation Transition of [41], separates a regime with a single energy minimum, from a disordered energy landscape with many minima. As discussed below, and demonstrated in Fig. 1, we find the same under-to-overparameterised/jamming transition, and show that an analogous TTT exists in the loss of DNNs as a function of the regularisation strength, $\lambda$.
@@
In order to obtain the phase diagram, we sample a uniform grid in the $(N,\lambda)$ plane, i.e. construct DNNs of varying size and regularisation, and train these networks with Stochastic Gradient Descent for 7000 epochs with a batch size of 100. We then perform 50 epochs of gradient descent at the original learning rate, and a further 50 epochs of gradient descent after reducing the learning rate by a factor of 10. Convergence was checked manually on a representative subset of DNNs. As such we are sampling (close to) the minima of the landscape. Finally we average over 10 DNNs with different, random initial conditions at each point in the $(N,\lambda)$ plane, resulting in the phase diagram shown in Fig. 1(a). We use SGD instead of Langevin dynamics as we are studying static properties of the landscape, not the dynamics, and SGD is much faster than Eq. (3). We include illustrative sketches of the resulting phases around the edge of Fig. 1(a) emphasising their important properties: the absence/presence of flat regions either side of the UOT, and the emergence of many minima below the TTT.
The UOT is visible on the $\lambda=0$ line, where at low $N/P_{D}$ the loss is approximately $10^{-2}$ (underparameterised), and at high $N/P_{D}$ falls to around $10^{-10}$ (overparameterised). There is some ambiguity as to where exactly the $\mathcal{L}=0$ phase appears, but we take the transition point to be $N^{\ast}/P_{D}\approx 0.18$. The TTT occurs in the vertical direction, where at high $\lambda$ the loss is 0.5, as is expected for a DNN where all weights are zero (see Eq. (1), with $f=0$). At low $\lambda$ the networks are able to find alternative solutions (i.e. with non-zero weights), and hence achieve losses of around 0.1. We take the TTT point to be where the loss falls below 0.5, hence $\lambda^{\ast}\approx 0.03$. In Appendix B, we demonstrate that the departure of the loss from a value of 0.5 corresponds to the landscape changing from having a single minimum at $w_{i}=0$ $\forall i$, to one with multiple distinct minima. The transitions are shown separately in Fig. 1(b) (UOT) and (c) (TTT). Figure 1(b) uses the same networks as on the $\lambda=0$ line from Fig. 1(a), and shows that achieving the $\mathcal{L}\approx 0$ state is by no means guaranteed. It is strongly dependent on the initial conditions and training trajectory through weight space, even in the overparameterised phase, in agreement with [35].
@@
On either side of the UOT we have two disordered phases, underparameterised, and overparameterised, illustrated schematically in Fig. 1(a). The key difference between these phases is the presence of flat plateaus at zero loss in the overparameterised phase. The UOT does not exist when $\lambda>0$ as the addition of the regularisation term destroys the plateaus. This is similar to the case where the addition of a spherical constraint makes a system of equations unsolvable [56]. Figure 1(c) uses 20 DNNs of width 30, each with different initial conditions, to demonstrate the TTT. Each DNN undergoes a sharp TTT, but the position of the transition varies, resulting in an average with some width. In this transition region a small number of initial conditions find a long-lived, high loss minimum, resulting in a kink in the average curve. The TTT separates a trivial phase, where the landscape is essentially a quadratic bowl (high $\lambda$), from a disordered phase (low $\lambda$).
@@
Figure 1: Glass-like transitions in neural networks. (a) The final training loss (colour bar) in the $(N/P_{D},\lambda)$ plane, where $N$ is the number of weights, $P_{D}$ the number of training data, and $\lambda$ the regularisation strength. Schematic illustrations of the loss landscape in the resulting four phases are shown around the edge. Note that these sketches are only illustrative and not based on real data. At high $\lambda$ the quadratic L2 regularisation term of Eq. (2) dominates and the loss is a simple well. As $\lambda$ decreases the loss landscape undergoes a TTT and develops many random minima. On the $\lambda=0$ line, at low $N/P_{D}$ the network is underparameterised, corresponding to another disordered landscape. Finally, at high $N/P_{D}$ the network is over parameterised and has a loss landscape combining disordered regions with flat plateaus at zero loss. (b) The final training loss either side of the UOT. Each black point is a different initial condition. For this plot $\lambda=0$. (c) The final training loss on either side of the TTT. The translucent lines correspond to different initial conditions, and the blue line is their average. For this plot $N/P_{D}=0.0805$, corresponding to the left most column in (a).
@@ III.2 Relaxation with overdamped Langevin dynamics
We now turn to the time-dependent dynamics of our DNNs to explore potential analogies with the relaxation dynamics of structural glassformers. There are four regions of interest in our phase diagram, on either side of each transition: the regularised phase ($0<\lambda<\lambda^{\ast}$); the quadratic phase ($\lambda>\lambda^{\ast}$); the underparameterised phase ($\lambda=0$, $N<N^{\ast}$); and the overparameterised phase ($\lambda=0$, $N>N^{\ast}$). In practice, the quadratic phase is trivial as networks rapidly reach steady state at the bottom of the quadratic well ($w_{i}\approx 0$ $\forall i$), hence we do not discuss this phase further. As can be seen from Fig. 1(a), $\lambda^{\ast}\approx 0.03$, with some $N$ dependence, and $N^{\ast}/P_{D}\approx 0.18$.
@@
We perform a set of numerical experiments designed to be qualitatively similar to the quench and anneal protocol used to study glasses (e.g. in [57, 58] and [59] among many others). Networks are first initialised with random weights drawn from a uniform distribution. Next, the networks are trained with SGD, at batch size 100, to efficiently move them away from their random initial position in the loss landscape, and into a local basin. Ideally this initial training period would be performed with Langevin dynamics, however SGD is significantly faster, and due to the long timescales involved in our measurements computing time was a limiting factor. The networks are then trained for a further period using the overdamped Langevin dynamics of Eq. (3), at a range of temperatures, $T$. After waiting some time, $t_{w}$, to allow the networks to move beyond their transient response to the new dynamics we start to measure $Q$. Measurements are made on approximately 50 independent initial conditions, although in some cases the network weights become unstable and grow to infinity. These networks are excluded from our analysis. A typical example of the measurement procedure is shown in Fig. 2(a). The decay of $Q$ at two temperatures is shown in Fig. 2(b). As expected, networks decorrelate more quickly at higher temperatures. We take as a typical decorrelation time $\tau_{90}$, the time it takes for $Q$ to reach 90% of its initial value, $Q(\tau_{90})=0.9$. Glassy systems typically undergo two step relaxation: an initial fast relaxation, followed by a caging plateau, and then slower relaxation in the long time limit. The $e$-fold relaxation time is used to probe the slower relaxation. However, we do not observe two step relaxation in any of our experiments, making the choice of overlap value somewhat arbitrary. As such, we choose $\tau_{90}$ for computational convenience as it is reached quickly. For completeness we repeat our measurements with $\tau_{e}$ and find no significant differences with $\tau_{90}$. These results are presented in Appendix E.
@@
Figure 2: (a) An example loss curve demonstrating the measurement procedure. Networks are generated with random weights, and trained for an initial period with SGD. The networks are then trained for a second period with overdamped Langevin dynamics. Measurements start after a waiting time, $t_{w}$. (b) Examples of the overlap function for networks in the high temperature, noise dominated regime (red, lower curve, $T=10^{-4}$), and in the low temperature, landscape dominated regime (blue, upper curve, $T=10^{-6}$). These measurements were made with $t_{w}=5\times 10^{3}$, and using $\chi=400$. The analytic predictions of $Q(t)$ for free Brownian particles are shown by the dashed black lines.
@@
In Fig. 3(a), (b), and (c), $Q(t)$ is plotted for a range of temperatures in the regularised (width=30, $\lambda=10^{-2}$), underparameterised (width=30, $\lambda=0$), and overparameterised (width=72, $\lambda=0$) phases respectively. The corresponding dependence of $\tau_{90}$ on the temperature, $T$, is shown in (d), (e), and (f). Both the regularised and underparameterised phases display two dynamical regimes. At high $T$, decorrelation is dominated by thermal noise, and we find a power law relationship, $\tau_{90}\sim T^{-1}$, which we call the noise-dominated regime. This is what we would expect from an overdamped, memoryless system with dynamics of the form $\dot{Q}=-\Omega Q$, where the frequency term $\Omega\sim T$ (e.g. Eq. 28 in [60] if we neglect the memory kernel). Interestingly, a $1/T$ relationship has also been observed for the decorrelation of the loss of DNNs trained on random data, and using the cross entropy loss function [61]. At low $T$, the decorrelation process becomes temperature independent, and instead is dominated by the first term of Eq. (3), the deterministic gradients of the loss landscape. The mean value of $\tau_{90}$ at $T=0$ is shown with the horizontal black line. Consequently, the value $T^{*}$ where plots (d), (e), and (f) exhibit a kink is the point at which the noise term of Eq. (3) can effectively be neglected. We hypothesise that this is because the networks become trapped in the basins of the loss landscape. We call this low $T$ behaviour the landscape-dominated regime. In the overparameterised phase only the noise-dominated regime, $\tau_{90}\sim T^{-1}$, is observed. We suggest that this is because the landscape consists of flat plateaus at $\mathcal{L}\approx 0$ where there are no gradients and hence the noise term of Eq. (3) always dominates. Furthermore, the observed divergence in Fig. 3(f) of $\tau_{90}$ as $T\to 0$ is to be expected, as both terms in Eq. (3) disappear.
Note the parameter $\epsilon$, that defines the overlap region in Eq. (5), is not the same between phases in Fig. 3, hence we cannot make quantitative comparisons between the measurements for the regularised, under-, and overparameterised networks. This stems from the lack of a typical scale for the weights. Choosing a fixed $\chi$ would also not result in comparable results as the standard deviation of the weights, $\sigma$, varies between phases. However, quantitative comparisons can be made between different temperatures in the same phase, as $\chi$ is fixed and $\sigma$ varies by around 1%. Furthermore, the key conclusions that two regimes exist (noise and landscape dominated), and that they behave as a $T^{-1}$ power law and a temperature independent regime, are robust to the changes in $\epsilon$ between phases. We cannot rule out that there exists some $\epsilon$ small enough that a landscape dominated regime emerges in the overparameterised phase. However, using the smallest $\epsilon$ possible without introducing issues with floating point precision (at approximately $\epsilon=10^{-7}$), we do not observe any deviation from the noise dominated regime.
@@
where $\text{erf}(x)$ is the error function. This analytic curve is shown in the dashed line of Fig. 2(b) where, for a network in the noise-dominated regime, the prediction closely follows the DNN measurement at short and intermediate times. In the landscape-dominated regime, the DNNs decorrelate faster than would be predicted by Eq.  (7) for Brownian weights at the equivalent temperature. Weights in a harmonic basin subjected to thermal noise decorrelate faster the steeper the basin (see Appendix F for details), hence, although we cannot perform analytic calculations with the true loss landscape, it is to be expected that DNNs in the landscape-dominated regime decorrelate faster than the prediction of Eq. (7).